# Accurate Low Rank Approximation at a Low Computational Cost

### v13

@@ Abstract
Given an algorithm $\mathbb{A}$ for computing a crude Low Rank Approximation (LRA) of a matrix having fast decaying spectra of singular values, we greatly decreased the spectral norm of its output error matrix at sublinear additional computational cost. Furthermore, for a variety of synthetic and real world inputs we tested numerically variants where our entire algorithms run at sublinear cost, and in most cases we still obtained accurate or even near-optimal LRAs. Towards enhancing our progress to a larger class of matrices, we nontrivially extended to LRA the popular technique of iterative refinement.
@@
low rank approximation, near-optimal algorithms, sublinear cost algorithms, singular values, gaps in the spectrum, iterative refinement
@@
1 LRA of a matrix: State of the Art and our Novelties
@@ 1.1 The State of the Art
Low rank approximation (LRA) of a matrix is a hot topic of Numerical Linear Algebra. [^1] The terms “low”, “large”, “small”, “close”, “near” etc are used in context in our introductory comments. Next we outline its State of the Art, referring for broader expositions to foremost papers [20, 54].
1.1.1. Some assumptions. For simplicity we seek LRA just in the field $\mathbb{R}$ of real numbers but can readily extend our studies to the complex field $\mathbb{C}$. We assume that an $m\times n$ input matrix $M$ is a black box matrix, given by an oracle (black box subroutine) for computing its product with a vector rather than by the entries. This is a natural model, e.g., for the study of matrices of immense sizes coming from Big Data, for which the access to all matrix entries is too costly (see further motivations in [54, Sec. 1.3]). Under this model linear operations with matrix-by-vector products can be performed with vectors.
Hereafter $a\ll b$ means that $a=o(b)$ but also that $b$ greatly exceeds $a$ in non-asymptotic context, $A^{+}$ denotes the Moore-Penrose pseudo inverse of $A$, and $m\geq n$; otherwise we can work with the $n\times m$ transpose $M^{T}$.
[…]
Computation of $M_{r}$ and $r$-top SVD involves order of $mn^{2}$ floating point operations (flops) and the memory space for about $mn$ scalars (see [15, Fig. 8.6.1]), although the output $r$-top SVD only uses memory for $(m+n+r)r$ scalars.
1.1.3. Variations and computational equivalence of LRA tasks. Next we state LRA task by generalizing this and other known representations of LRA as follows: given a matrix $M\in\mathbb{R}^{m\times n}$ and a pair of integers $r_{-}$ and $r_{+}$ such that
$0<r_{-}\leq r_{+}\ll n\leq m,$ (1.1)
fix two integers $k$ and $\ell$ such that
$k\ll m,~{}\ell\ll n,~{}r\leq\min\{k,\ell\},~{}{\rm and}~{}r_{-}\leq r\leq r_{+}$ (1.2)
and seek a rank-$r$ approximation of $M$ by a triplet of matrices $X\in\mathbb{R}^{m\times k}$, $Y\in\mathbb{R}^{k\times\ell}$, and $Z\in\mathbb{R}^{\ell\times n}$ such that the error matrix $M-XYZ$ has a small spectral and/or Frobenius norm. [^3] For $r_{-}=r_{+}$, relationships (1.2) define a target rank $r=r_{-}=r_{+}$. Otherwise, more realistically, they define a target range for rank $r$.
If the matrices $X=C$ and $Z=R$ are made up of the columns and rows of $M$, respectively, then $Y$ is commonly denoted by $U$ (or by $G$ by the followers of the pioneering work [17]) and the triplet $\{X,Y,Z\}$ is called CUR LRA (or CGR LRA) and is a particularly memory efficient LRA.
For any $r\leq\min\{k,\ell\}$ we can transform any LRA triplet $\{X,Y,Z\}$ for $M$ into its $r$-top SVD and further into its CUR (CGR) LRA by using $O(mk^{2}+n\ell^{2})$ flops, $(m+k)k+(n+\ell)\ell$ scalars and $O(mr^{2})$ flops, $(m+n+r)r$ scalars, respectively (see Appendices B and C.2, respectively).
1.1.4. LRA via random sampling. Impressive progress in LRA in the last decades relied on random sampling algorithms (see [8, 6, 26, 20, 56], and the references therein).
We need an efficient LRA algorithm for incorporation into our Alg. 2.1 of Sec. 2 and for natural candidates consider random sampling algorithms, such as the following algorithm of [54], which modifies those of [6, Thm. 4.7] and [6, Thm. 4.3, display 1].
Algorithm 1.1.
[LRA by means of random sampling].
Input:
An $m\times n$ black box matrix $M$, given implicitly by a subroutine for its multiplication by vectors, and positive integers $r_{-}$ and $r_{+}$ satisfying (1.1).
Output:
An integer $r$ satisfying (1.2) and, moreover, equal to $\min\{k,\ell\}$ and a triplet of matrices $\{X,Y,Z\}$ such that $XYZ$ is a rank-$r$ matrix close to $M$.
Initialization:
Fix a pair of integers $k$ and $\ell$ satisfying (1.2) and generate a pair of independent random test matrices $F\in\mathbb{R}^{\ell\times m}$ and $H\in\mathbb{R}^{n\times k}$.
Computations:
  1. 1.
    Compute the matrices $Z:=FM\in\mathbb{R}^{\ell\times n}$ and $MH\in\mathbb{R}^{m\times k}$, which are two sketches (samples) of $M$.
  2. 2.
    Compute the matrix $X\in\mathbb{R}^{m\times k}$ denoting the Q factor of the thin QR factorization of $MH$ – (cf. [15, Thm. 5.3.2]).
  3. 3.
    Compute the matrx $Y:=(FQ)^{+}\in\mathbb{R}^{k\times\ell}$.
  4. 4.
    Output the triplet of matrices $\{X,Y,Z\}$.
The algorithm involves $O((k+\ell)mn)$ flops [^4] One can do with just $O(mn\log(k+\ell))$ flops to ensure a little weaker bounds on failure probability [20, Sec. 11.2]. and about $(m+n)(k+\ell)$ scalars, not counting the cost of the generation of random matrices $F$ and $H$. Like various other random sampling LRA algorithms, Alg. 1.1 is universal, that is, devised for all matrices $M\in\mathbb{R}^{m\times n}$; for proper pairs of dense test matrices $f$ and $H$ and integers $k$ and $\ell$ in $O(r\log(r))$ it belongs to the class of relative error LRA algorithms: for any fixed pair of positive $\epsilon$ and $\delta$ it outputs rank-$r$ approximations of $M$ whose Frobenius error norm is within a factor of $1+\epsilon$ from the optimal value $||M-M_{r}||_{F}$ with a failure probability at most $\delta$ (see the bibliography cited above).
1.1.5. Superfast (sublinear cost) LRA. The LRA algorithms of [28, 4, 21, 23, 5] (randomized) and [24, Part III] (deterministic) are known to run superfast, aka at sublinear cost, that is, to use $o(mn)$ scalars and flops for a fixed pair of $\epsilon$ and $\delta$ [^5] Superfast algorithms have been studied extensively for Toeplitz, Hankel, Vandermonde, Cauchy, and other structured matrices having small displacement rank [34, 36], and we extend this concept to other matrices.
@@
Unfortunately, already for an input family of $mn+1$ matrices having size $m\times n$ and rank at most 1, any superfast algorithm at best halves the large error of the trivial approximation by the matrix filled with 0s (see Appendix A or, say, [40, 41, 42, 43]). If randomization is applied, such an algorithm still at best halves the error of the trivial approximation with a probability not close to 0. Likewise, any deterministic or randomized superfast estimation of the error norm of an LRA fails on the matrices of that small family. Hence a superfast algorithm cannot be universal for computation of LRA or even for its a posteriori error norm estimation.
Accordingly, the randomized LRA algorithms known to be superfast are restricted to some large and important classes of input matrices such as Symmetric Positive Semidefinite (SPSD) matrices, Distance matrices, or a subclass of Cauchy matrices. Furthermore, strictly speaking they are NOT superfast because they involve a posteriori error norm estimation in linear or superlinear time. The latter deficiency is only avoided in the deterministic algorithm of [24, Part III] for LRA of SPSD matrices – its output error norm is estimated as by-product, at no additional cost.
1.1.6. Conflicts #$\#1$ for LRA. Trying to optimize simultaneously various permissible parameters of LRA such as arithmetic cost and the costs of randomization, storage, and communication, representing the computational cost one runs into the conflict with the quality of a priori error bounds one (see [54, Sec. 3.9]), which boils down to the conflict between the two main goals of devising efficient LRA algorithms, that is, minimization of (a) the output error norm and (b) represented by the memory space and number of flops involved. Here are some specific samples of this omnipresent LRA conflict #$\#1$, routinely resolved by means of trade-off.
(i) The known random sampling LRA algorithms need unbounded memory size and flop number to ensure with a high probability (whp) the relative error norm bound within $1+\epsilon$ from optimal under the Frobenius norm for $\epsilon\rightarrow 0$ [20, Sec. 10.3]. Furthermore, the LRA algorithms supporting the same output error bound under the spectral rather than Frobenius norm are more costly (see [56, Ch. 6.2], [13, Thm. 3.4]).
(ii) Adaptive Cross-Approximation (ACA) iterations, [^6] We briefly recall their simplified version of Cross-Approximation (C-A) iterations in Appendix E. which are empirically superfast for a large class of matrices $M$, output LPAs whose error norm bound is at best within a factor of $2\sqrt{mn}$ from optimal under both spectral and Frobenius matrix norms (see [30, Cor. 2]).
(iii) Likewise, the output error norms of the relative error LRAs of [24, Part III] for $n\times n$ Symmetric Positive Semidefinite (SPSD) matrices $M$ may exceed the optimal spectral and Frobenius norms $||M-M_{r}||$ by a factor of $n$.
1.2 Our progress (briefly)
For matrices with fast decaying spectra of singular values (hereafter FDS matrices) our algorithms decrease the spectral norm of the output error matrix of the known LRA algorithms at sublinear additional cost, that is, superfast. Our numerical tests with both synthetic and real world matrices are in good accordance with these formal estimates. Furthermore, some variants of our entire algorithms run superfast, and we empirically observe that they still output accurate or even near-optimal LRAs for a large subclass of FDS matrices.
1.3 LRA by means of escalation
Next we outline our rather simple techniques.
1.3.1. More definitions and two theorems. Hereafter $\sigma_{j}(M)$ denotes the $j$th top (largest) singular value of an $m\times n$ matrix $M$, $j=1,2,\dots$; $\sigma_{j}(M)=0$ for $j>\operatorname{rank}(M)$. $||X||:=\sigma_{1}(X)$ denotes the spectral norm of a matrix $X$, to which we restrict our study, except for Sec. 5.
In particular we only recall the Eckart-Young-Mirsky theorem (cf. [50, 12, 27], [15, Thm. 2.4.8]) under the spectral norm:

### v16

# Near-Optimal Very Low Rank Approximation at Sublinear Cost

@@ Abstract
Eckart-Young’s theorem implies that even optimal rank-$r$ approximation of an $m\times n$ matrix $M$ for $m\geq n$ is only interesting if there is a gap in the segment of the $r+1$ largest singular values of $M$. If $M$ also has a gap in the segment of the next $p$ largest singular values, then one can compute its rank-$r$ approximation with random sampling and oversampling in nearly linear arithmetic time. For matrices $M$ having such a double gap in their spectra and for $n$ greatly exceeding $(r+p)^{2}$, our algorithm runs in sublinear time, and for a large subclass of the class of such matrices $M$ its output rank-$r$ approximation of $M$ is near-optimal under the spectral norm. We also extend the classical iterative refinement to LRA and analyze and test the resulting algorithm.
@@
low rank approximation, singular values, gaps in the spectrum, sublinear cost algorithms, iterative refinement
@@
1 LRA of a matrix: the State of the Art and our Novelties
Low rank approximation (LRA) of a matrix is a hot topic of Numerical Linear Algebra. We first briefly recall the State of the Art, referring to [11, 33, 18] for broader expositions, and then outline our LRA algorithms.
@@ 1.1 The State of the Art
1.1.1. LRA and CUR LRA. Assume that a black box matrix $M\in\mathbb{C}^{m\times n}$ is given by an oracle (black box subroutine) for computing its product with a vector rather than by its entries; then linear operations with matrices can be reduced to linear operations with vectors. [^1] The sketch of a matrix in [36, 33] is its product with a fixed set of vectors. Then we can handle matrices of immense size that come from the study of Big Data and are too large to be accessed otherwise (see further motivations in [33, Sec. 1.3]). Furthermore, quite typically the matrices representing Big Data are close to low rank matrices [35], that is, admit LRA, with which one can operate superfast, aka at sublinear cost – by using much fewer than $mn$ scalars and flops. [^2] Superfast algorithms have been studied extensively for Toeplitz, Hankel, Vandermonde, Cauchy, and other structured matrices having small displacement rank [24], and we extend this concept to other matrices.
Now assume that $m\geq n$ (otherwise we could have worked with the $n\times m$ transpose $M^{T}$) and that we are given a positive target rank $r\ll n$ or a fixed target range [^3] Here and hereafter $a\ll b$ means that $b$ greatly exceeds $a$ in context, e.g., $a=o(b)$ as $b\mapsto\infty$. for such a rank (actually, a proper choice of such targets is a challenge [11, Sec. 4.2]) and then define LRA of $M$ by a triplet $\{X,Y,Z\}$ such that
$X\in\mathbb{C}^{m\times k},Y\in\mathbb{C}^{k\times\ell},~{}Z\in\mathbb{C}^{\ell\times n},$ (1.1)
$k\ll m,~{}\ell\ll n,~{}\min\{k,\ell\}=r,$ (1.2)
and the spectral and/or Frobenius norm of the matrix $M-XYZ$ is small in context, say, close to optimal.
The matrices $X,Y,Z$ have $mk+k\ell+\ell n\ll mn$ entries overall, and we can multiply the matrix $XYZ$ by a vector by using less than $2(mk+k\ell+\ell n)\ll mn$ flops.
If the matrices $X$ and $Z$ are made up of the columns and rows of $M$, that is, are defined by means of pre- and post-multiplication of $M$ by selected coordinate vectors, respectively, then write $X:=C$, $Z:=R$, and $Y:=U$ and call the triplet $\{X,Y,Z\}$ a CUR LRA.
1.1.2. LRA from top SVD, SVD of an LRA, and CUR from SVD. By setting to 0 all singular values of $M$ except for its $r$ top (largest) ones, obtain a matrix $M_{r}$ of rank $r$, said to be $r$-truncation of $M$. By virtue of the Eckart-Young-Mirsky theorem [7, Thm. 2.4.8], $M_{r}$ is an optimal rank-$r$ approximation444“Rank- $r$ approximation” means “rank at most $r$ approximation” throughout. of $M$ under both spectral and Frobenius norms, that is, $X=M_{r}$ minimizes these norms of $M-X$ over all matrices $X$ of rank $r$.
Computation of SVD of $M$ involves memory space for more than $mn$ scalars and order of $mn^{2}$ floating point operations (flops) (see [7, Fig. 8.6.1]). This is too expensive for computing LRA of a large size matrix, while conversely, given a triplet $\{X,Y,Z\}$ satisfying (1.1) and (1.2), we can compute SVD of $XYZ$ and then its CUR LRA by using $(m+k)k+(n+\ell)\ell$ scalars, $O(mk^{2}+n\ell^{2})$ flops and $(m+n+r)r$ scalars, $O(mr^{2})$ flops, respectively (see Appendices B and C.2, respectively).
1.1.3. LRA via random sampling and oversampling. Random sampling with oversampling enables fast computation of LRA (see [3, 17, 11, 36, 33, 20], and the references therein). For any $M\in\mathbb{C}^{m\times n}$ and target rank $k<n\leq m$, random sampling HMT algorithm of [11] successively generates a random test matrix $H\in\mathbb{C}^{n\times(k+p)}$ for some positive oversampling integer $p$, usually a small constant, computes the product $MH$, and computes rank-$k$ approximation of $M$.
Remark 1.1.
“In practice, the target rank $k$ is rarely known in advance. Randomized algorithms are usually implemented in anadaptive fashion where the number of samples is increased until the error norm satisfies the desired tolerance” (see [11, Sec. 4.2]).
In the best studied case the test matrix $H$ is random Gaussian, filled with independent standard Gaussian (normal) random variables. In that case an expected Frobenius error norm of the output LRA is within a factor of $\epsilon=(1+\frac{k}{p-1})^{1/2}$ from the optimal value $||M-M_{r}||_{F}$ [11, Thm. 10.5], and the algorithm uses orders of $(m+n)(k+p)$ scalars and of $(k+p)mn$ flops. The parameter $p$ and the size of the test matrix $H$ explode as $\epsilon\mapsto 0$, but LRA with $\epsilon$ as large as 1/2 or 1 is still valuable.
In other LRA algorithms (e.g., in [4, 3, 33, 20]) $M$ is both pre- and post-multiplied by random test matrices $F$ and $H$, respectively. In particular, the TYUC algorithm of [33], modifying those of [3, Thm. 4.7] and [3, Thm. 4.3, display 1], and the generalized Nyström approximation GN of [20] first compute the products $FM$ and $MH$ for random test matrices $F\in\mathbb{C}^{k\times m}$ and $H\in\mathbb{C}^{n\times\ell}$ for $k\leq\ell$, say, for $\ell=2k$, and then compute rank-$k$ approximation of $M$ with the expected Frobenius error norm within a factor of 2 from optimal in the case of random Gaussian test matrices $F$ and $H$ (namely, the norm does not exceed the optimal one for rank-$\frac{k}{2}$ approximation). For $\ell=2k$ the TYUC and GN algorithms use about $(m+n)\ell$ scalars (or slightly less in [20]) and $O(\ell mn)$ flops.
The computation becomes a little more costly for supporting the same LRA error bounds under the spectral rather than Frobenius norm (see [11, Secs. 10 and 11], [36, Ch. 6.2], [5, Thm. 3.4]).
1.1.4. Superfast LRA for special subclasses of matrices. For some special but large and important classes of matrices, most notably for Symmetric Positive Semidefinite (SPSD) matrices but also for Distance matrices and a subclass of Cauchy matrices, the superfast randomized algorithms of [19, 1, 12, 14, 2] compute LRA whose expected Frobenius error norm is compared well with the optimal bound. The mere existence of such LRA algorithms is surprising for $n\times n$ SPSD matrices, whose class includes matrices $A^{*}A$, defined by nonsingular $n\times n$ matrices $A$ with $n^{2}$ independent entries. Notice that the superfast algorithm of [15, Part III] for SPSD matrices is deterministic, but its spectral and Frobenius output error norms exceed their optimal values by at most a factor of $n$.
@@
Unfortunately, one cannot compute LRA superfast for every matrix: already for an input family of $mn+1$ matrices having size $m\times n$ and rank at most 1, any superfast algorithm at best halves the large error of the trivial approximation by the matrix filled with 0s (see Appendix A or, say, [26, 27, 28]). If randomization is applied, such an algorithm still at best halves the error of the trivial approximation with a probability not close to 0.
1.2 Our LRA algorithm
Hereafter $\sigma_{j}(M)$ denotes the $j$th largest singular value of an $m\times n$ matrix $M$, $j=1,2,\dots$; $\sigma_{j}(M)=0$ for $j>\operatorname{rank}(M)$, and let $M$ be an $(r,\rho)$-matrix, that is, let it have gaps in both segments $[\sigma_{1}(M),\sigma_{r+1}(M)]$ and $[\sigma_{r+1}(M),\sigma_{\rho+1}(M)]$ of its singular values for $r<\rho\leq n\leq m$. We seek rank-$r$ approximation of $(r,\rho)$-matrices $M$ of size $m\times n$ such that $r<\rho\ll\sqrt{n}$. Our algorithm is superfast and, for a large subclass of $(r,\rho)$-matrices $M$, its output very low rank approximation (VLRA) is near-optimal under spectral norm.
We rely on combining three surprisingly simple observations, namely, that (i) random sampling with oversampling is the only stage where HMT, TYUC, and GN algorithms are not superfast for VLRA; (ii) we can perform this stage superfast by using sparse test matrices (in particular abridged SRHT matrices of Appendix D), and (iii) while the accuracy of the VLRA output in step (ii) tends to be unsatisfactory, we can readily improve it to a near-optimal level for a large subclass of $(r,\rho)$-matrices $M$. Our LRA Alg. 2.1 of Sec. 2 relies on these observations.
We compute at first (1) a crude but reasonably close rank-$\rho$ approximation $M(\rho)$ of $M$ and then (2) the rank-$r$ truncation $M(\rho)_{r}$ of $M(\rho)$, which we output as rank-$r$ approximation of $M$.
LRA via random sampling with oversampling fits this outline, but based on above observations (i)–(iii) we run the algorithm superfast for $\rho\ll\sqrt{n}$ and argue that its output VLRA is near-optimal for a large subclass of $(r,\rho)$-matrices.
We can perform stage (1) of the algorithm superfast by applying any of the HMT, TYUC and GN algorithms for $k:=\rho\ll\sqrt{n}$ with sparse test matrices such that their multiplication by $M$ is superfast (see [13], [33, Sec. 3.9], and our Sec. 3).
Apart from the stage of multiplication by test matrices, the algorithm HMT uses $(k+p)m$ scalars and $O((k+p)^{2}m)$ flops, while the TYUC and GN algorithms use about $mk+n\ell$ and $mk+n\ell+1.5kl$ scalars, respectively, and $O((m+k)k\ell)$ flops, for $k\leq\ell$, and obtain Moore-Penrose pseudoinverses of a $k\times\ell$ matrices ${\bf\Psi Q}$ in [33] and $R_{\epsilon}$ in [20] by using $O(k^{2}\ell)$ flops (see [7, Fig. 8.6.1]). These computations are superfast where $(k+p)^{2}\ll n$ and $k^{2}\ll n$, respectively, for $\ell\leq 2k$, say.
At Stage (2) we apply Alg. B.1. It computes SVD and $r$-truncation of $M(\rho)$ by using about $(m+n)\rho$ scalars and $O(mr^{2})$ flops, that is, $\ll m\sqrt{n}$ scalars and $\ll mn$ flops for $\rho\ll\sqrt{n}$.
By virtue of Thm. 2.3, the spectral norm $||\cdot||:=||\cdot||_{2}$ of the output error matrix of Alg. 2.1 satisfies
$||M-M(\rho)_{r}||\leq||M-M_{r}||+2||M-M(\rho)||,$ (1.3)
that is, Alg. 2.1 is optimal up to a factor of $1+2\frac{||M-M(\rho)||}{||M-M_{r}||}$. For $(r,\rho)$-matrices, this factor can be large only if the superfast computation of a reasonably close rank-$\rho$ approximation of $M$ at stage 1 fails. It does fail for worst case matrices $M$ (see Appendix A); moreover, empirically the output matrices of the algorithms HMT, TYUC, and GN with sparse test matrices tend to be unsatisfactory as LRA [13], [33, Sec. 3.9], but in our numerical tests in Sec. 5 accuracy of the output matrices of stage (1) of Alg. 2.1 tended to be raised to near-optimal level at stage (2).
1.3 Iterative LRA
In Sec. 4 we apply our LRA algorithm to the output error matrix, hoping to decrease its norm if it is too large. Recursively we extend to LRA the popular techniques of iterative refinement and approximate $M$ with a sum of computed LRAs. The algorithm is essentially reduced to recursive applications of Alg. 2.1, but the rank of computed sums and hence the cost of their computation grow fast as recursive steps advance. We counter this deficiency by recursively compressing the sums into rank-$r$ matrices, although this limits output accuracy of computed LRAs. The resulting iterative LRA algorithm is a natural extension to LRA of an efficient popular technique of matrix computations; we specify, analyze, and test it in Secs. 4 and 5.
We devised this algorithm to complement and strengthen Alg. 2.1, but it is essentially reduced to recursive applications of Alg. 2.1. Our upper estimates for the accuracy of both algorithms are close to one another, although both can be quite pessimistic (see Remarks 2.1 and 4.2).
In our tests, the output accuracy of our iterative LRA applied to $(r,\rho)$-matrices improved significantly at the second iteration, little improved afterwards, and overall error norm was never less than with Alg. 2.1 for the upper rank $\rho=4r$.
1.4 Related works
Our study was prompted by extensive work on LRA in the last decades (see [11, 33, 18] and the bibliography therein) and in particular by fast LRA based on random sampling and oversampling. To accelerate these algorithms we use abridged SRHT matrices studied in the papers [26, 27, 28], which in turn extends the earlier study of Gaussian elimination without pivoting in [29, 31].
Our iterative LRA extends the algorithms of [16, 25], which in turn extend to LRA the classical technique of iterative refinement, particularly popular for the solution of a linear system of equations (see [32, Secs. 3.3.4 and 4.2.5], [10, Ch. 12 and Sec. 20.5], [7, Secs. 3.5.3 and 5.3.8]) but also widely applied in various other linear and nonlinear matrix computations (see [32, page 223 and the references on page 225]). Its extension to LRA is not straightforward because one must apply re-compression to block recursive increase of the rank of the computed approximations; in a sense we extend re-compression of [22, 30].
Actually, we first devised iterative refinement of LRA that incorporated our oversampling technique and then noticed that that technique alone can be competitive or superior.
1.5 Organization of our paper
We present our accelerated oversampling LRA algorithm in Sec. 2, discuss superfast computation of a crude LRA in Sec. 3, present our iterative LRA algorithm in Sec. 4, and cover our numerical experiments in Sec. 5. We devote our short Sec. 6 to conclusions. In Appendix A we describe some small families of matrices for which any superfast LRA algorithm fails. In Appendix B we compute $r$-top SVD of an LRA of a low-rank matrix superfast. In Appendix C we recall CUR decomposition and CUR LRA and transform an $r$-top SVD of a matrix into its rank-$r$ CUR LRA. In Appendix D we generate abridged SRHT matrices, whose incorporation as test matrices makes random sampling LRA superfast.
2 LRAs of $(r,\rho)$-matrices: generic algorithm and its analysis
We begin with some background materials.
Hereafter $||X||:=\sigma_{1}(X)$ denotes the spectral norm of a matrix $X$, to which we restrict our study, except for Remark 2.2. In particular we only use Eckart-Young’s (rather than Mirski’s) theorem: